### v2

# Think Outside the Policy: In-Context Steered Policy Optimization

Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts which are confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advanced models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces mixed-policy GRPO with implicit expert forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates expert region reject sampling to filter unreliable off-policy trajectories and annealed expert-bonus reward shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances RLVR performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs. Our code is available at https://anonymous.4open.science/r/ICPO.
Think Outside the Policy: In-Context Steered Policy Optimization
Hsiu-Yuan Huang1,2thanks:   Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3thanks:   Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1, [^2] footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉  huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
1 Introduction
Figure 1: Illustration of optimization dynamics in parameter space. GRPO exploration is confined to the current policy’s distribution, limiting trajectory diversity and often leading to suboptimal convergence. While prior methods expand exploration by incorporating expert rollouts generated by stronger LRMs on the training data, ICPO leverages existing datasets as in-context guidance, eliminating reliance on advanced LRMs.
Large Reasoning Models (LRMs) excel at solving complex mathematical problems, and Reinforcement Learning from Verifiable Rewards (RLVR) provides a scalable way to refine their reasoning through verifiable rewards DeepSeek-AI (2025). Yet, limited exploration under standard Group Relative Policy Optimization (GRPO) Shao et al. (2024) often hinders robust reasoning improvement Yue et al. (2025); Zhang et al. (2025a).
Recent work has explored combining Supervised Fine-Tuning (SFT) with Reinforcement Learning (RL) to strengthen the exploration of LRMs. One research direction interleaves SFT and RL updates Ma et al. (2025), enabling SFT to improve high-difficulty problem-solving while RL refines mid- and low-difficulty behaviors. However, repeated switching between paradigms introduces instability and inefficient convergence. Another line of work seeks to unify SFT and RL within a single training process through different strategies: incorporating SFT data as off-policy rollouts during RL to expand the exploration space Yan et al. (2025); jointly optimizing SFT and RL objectives for tighter integration Fu et al. (2025); and leveraging hints to bootstrap rollouts to improve performance on harder prompts Liu et al. (2025a); Zhang et al. (2025b); Fu et al. (2025).
Together, these approaches reflect a growing consensus that combining SFT and RL can effectively expand the exploration of policy search into more promising reasoning spaces. However, as shown in Figure 1, three key challenges remain: (1) Online exploration remains confined to the current policy distribution, as GRPO-based methods rely on on-policy sampling, resulting in limited trajectory diversity and might converge to local optima. (2) Expanding the exploration space with trajectories from stronger LRMs incurs high computational cost and limited accessibility, since generating additional reasoning traces from advanced models is expensive and such models are not always available. (3) External trajectories are often noisy and unstable for training, incorporating them indiscriminately as off-policy rollouts can mislead policy updates and harm convergence stability.
To address these limitations, we propose In-Context Steered Policy Optimization (ICPO), a novel RLVR framework that exploits the LRM’s inherent In-Context Learning (ICL) capability to provide expert guidance instead of relying on external advanced LRMs. Specifically, (1) ICPO introduces mixed-policy GRPO with Implicit Expert Forcing (IEF), where expert-conditioned rollouts are generated through ICL guidance, enabling exploration beyond the current policy distribution and steering the model toward expert-aligned regions of the solution space. (2) To ensure reliable guidance, ICPO employs Expert Region Reject Sampling (ERRS), which filters out noisy or low-quality off-policy trajectories using verifiable reward signals, preventing misleading gradients from contaminating policy updates. (3) ICPO further explores an annealed expert bonus into the Reward Shaping (RS) design, enforcing strong expert-guided shaping in the early stage and progressively relaxing it to facilitate autonomous optimization as LRM capabilities grow.
Our experiments show that ICL-based rollouts exhibit superior quality and diversity, making them effective sources of expert guidance (§ 3). Extensive evaluations across multiple mathematical reasoning benchmarks further demonstrate that ICPO achieves substantial performance gains over vanilla GRPO and mixed-policy GRPO Yan et al. (2025), with maximum average improvements of up to +4.1 and +4.0 points, respectively (§ 6). In addition, we provide in-depth instance- and token-level analyses that offer strong evidence of distributional shifts induced by ICPO (§ 7).
Our contributions are three-fold:
  • We empirically show that ICL rollouts provide diverse and high-quality expert signals for the mathematical reasoning task.
  • We introduce ICPO, which includes mixed-policy GRPO with Implicit Expert Forcing, leveraging the model’s inherent ICL ability without relying on stronger LRMs, together with Expert Region Reject Sampling and Reward Shaping for stable and efficient learning.
  • We empirically validate that ICPO delivers consistent improvements on mathematical reasoning benchmarks across model scales and effectively encourages exploration, demonstrating strong potential for LRM post-training.
2 Related Work
Exploration from Within the Policy.
Recent methods enhance exploration by increasing diversity within the policy itself, such as generating additional rollouts Wang et al. (2025); Hu et al. (2025a); Tang et al. (2025), leveraging replay buffers to revisit informative prompts Dou et al. (2025), and sampling under higher temperatures to boost stochasticity Chen et al. (2025); Zhang et al. (2024). These approaches broaden trajectory coverage under the model’s own distribution. However, their exploration remains bounded by the intrinsic capacity of the current policy.
Exploration from Outside the Policy.
Other studies leverage external supervision to guide exploration beyond the model’s default policy distribution. One line of work explores hybrid SFT–RL strategies to expand the reasoning space of LRMs. ReLIFT Ma et al. (2025) alternates between RL and SFT by updating on failed rollouts. LUFFY Yan et al. (2025) incorporates SFT trajectories as off-policy samples using importance sampling. SRFT Fu et al. (2025) jointly optimizes SFT and RL objectives with an entropy-based weight on the SFT loss. Another line of work guides rollouts by concatenating partial SFT solutions as hints Liu et al. (2025a); Zhang et al. (2025b); Huang et al. (2025). However, these methods rely on advanced LRMs to supply SFT traces, which may incur additional computation overhead.
In contrast, we exploit LRMs’ inherent ICL ability to steer diverse rollouts with existing datasets as demonstrations, requiring neither external LRMs’ trajectories nor explicitly engineered hints.
3 Preliminary
Explicit Expert Forcing.
In traditional RL and RLHF, expert forcing explicitly constrains the policy to align with an expert policy $\pi_{\phi}$, typically through imitation or KL-based regularization Hester et al. (2018); Haldar et al. (2023); Zhang et al. (2023); Hu et al. (2023). This explicit constraint stabilizes optimization and reduces reward variance, but it requires gradient-based imitation and access to an auxiliary expert model (e.g., a larger LRM), which can limit exploration in later stages.
ICL as Expert-Conditioned Inference.
ICL provides a gradient-free way to inject expert priors through the input context, which we evaluate from three complementary perspectives. (1) Accuracy: the 1-shot ICL setting consistently outperforms the 0-shot baseline (Figure 2), indicating that conditioning on demonstrations improves reasoning correctness. (2) Diversity: compared with temperature-based sampling, introducing 1-shot demonstrations expands the sampling space (Figure 3), yielding larger inter-trajectory edit distances and enhanced exploratory diversity. (3) Distribution quality: under ICL-conditioned rollouts, the output distribution becomes more favorable—a higher proportion of previously incorrect generations are “flipped” to correct solutions compared with temperature perturbations (Figure 3), indicating that in-context steering provides a stronger and more targeted exploration signal. Taken together, these results support our view that ICL constitutes an effective expert-conditioned inference process.
From ICL to Implicit Expert Forcing.
Given expert demonstrations $\mathcal{D}$ and a query $q$, the model generates trajectories conditioned on:
$x_{\mathrm{exp}}=[\mathcal{D};q],\quad\tau_{\mathrm{exp}}\sim\pi_{\theta}(\tau\mid x_{\mathrm{exp}}).$ (1)
Following the hypothesis-class view of ICL Hendel et al. (2023), the forward process of a Transformer $T$ can be decomposed into two functions:
$T([\mathcal{D},q])=\mathcal{F}(q;A(\mathcal{D})),$ (2)
where $A(\cdot)$ maps demonstrations $\mathcal{D}$ to a task vector $\vartheta=A(\mathcal{D})$ that encodes the expert behavior specific to that task Hendel et al. (2023); Li et al. (2023); Todd et al. (2024); Huang et al. (2024); Liu et al. (2024); Hojel et al. (2024), and $\mathcal{F}(\cdot\,;\vartheta)$ represents the task-specific reasoning function that applies the $\vartheta$ to generate the prediction for query $q$. This leads to the parametric representation:
$\pi_{\theta}^{\text{IEF}}(\tau\mid q):=\pi_{\theta}(\tau\mid[\mathcal{D};q])=\pi_{\mathcal{F}}(\tau\mid q;\vartheta),$ (3)
indicating that ICL implicitly introduces an expert-induced prior $\vartheta$ that steers the rollout distribution toward expert-like regions—without any explicit optimization on $\pi_{\theta}$. While ICL itself is an inference-time mechanism that does not update model parameters, we incorporate its induced trajectories into GRPO training to form IEF.
Figure 2: Comparison between 0-shot and 1-shot ICL on reasoning accuracy across benchmark datasets.
Figure 3: Effect of in-context steering on exploration and diversity. Temp. increases the default decoding temperature (0.6→1.2), and Expert. introduces in-context guidance (0-shot→1-shot). Red dots indicate instances flipped from incorrect to correct.
Group Relative Policy Optimization (GRPO).
GRPO is an efficient On-Policy optimization algorithm tailored for RL in LLMs, where the advantage for each token is computed in a group-relative manner without requiring an additional critic model to estimate token values. Given a set of rollouts $\{\tau_{i}\}_{i=1}^{N}$ sampled from the old policy $\pi_{\theta_{\text{old}}}$, the normalized advantage is computed by:
$A_{i}=\frac{R(\tau_{i})-\mathrm{mean}(G)}{\mathrm{std}(G)},\quad G=\{R(\tau_{i})\}_{i=1}^{N}.$ (4)
Analogous to PPO Schulman et al. (2017), the GRPO objective is formulated as:
$\mathcal{J}_{\scriptscriptstyle\text{GRPO}}(\theta)=\frac{1}{\sum_{i=1}^{N}|\tau_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|\tau_{i}|}\text{CLIP}(r_{i,t}(\theta),A_{i},\epsilon),$ (5)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(\tau_{i,t}|\tau_{i,<t})}{\pi_{\theta_{\text{old}}}(\tau_{i,t}|\tau_{i,<t})}$ is the importance ratio, and $\text{CLIP}(r,A,\epsilon)=\min(r\cdot A,\text{clip}(r;1-\epsilon,1+\epsilon)\cdot A)$ is the clipping function for variance reduction. To prevent the learned policy from drifting too far from the reference model, we retain the KL regularization term $\beta\cdot D_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]$ in GRPO, which is jointly optimized to ensure training stability and maintain controllable policy updates.
By leveraging ICL-conditioned rollouts within a mixed-policy GRPO framework, our approach enables expert-guided exploration to directly participate in policy optimization, effectively realizing an In-Context Steered Policy Optimization process.
4 Method
Figure 4 illustrates the overall ICPO training framework and the ICPO training process is detailed in Algorithm 1.
Figure 4: Overall Framework of ICPO. ICPO performs mixed-policy GRPO using off-policy trajectories generated by the policy model itself via implicit expert forcing, with reject sampling and reward shaping to stabilize training.
4.1 Mixed-Policy GRPO with Implicit Expert Forcing
To incorporate expert-conditioned exploration into group rollouts, we follow Yan et al. (2025) and extend GRPO into a mixed-policy setting, where each group consists of $N_{\text{on}}$ on-policy trajectories $\tau_{i}\sim\pi_{\theta_{\text{old}}}$ and $N_{\text{off}}$ trajectories generate under IEF $\tau_{j}\sim\pi_{\theta_{\text{old}}}^{\text{IEF}}$, such that $N_{\text{on}}+N_{\text{off}}=N$. We can recompute the group-normalized advantage (as in Eq. 4) over the mixed rollout set as:
$\hat{A}_{i}=\frac{R(\tau_{i})-\text{mean}(G_{\text{on}}\cup G_{\text{off}})}{\text{std}(G_{\text{on}}\cup G_{\text{off}})},$ (6)
where $G_{\text{on}}=\{R(\tau_{i})\}_{i=1}^{N_{\text{on}}}$ and $G_{\text{off}}=\{R(\tau_{j})\}_{j=1}^{N_{\text{off}}}$.
The objective of mixed-policy GRPO with IEF balances exploitation of the current policy with exploration toward expert-aligned regions, and can be written as follows:
$\mathcal{J}_{\scriptscriptstyle\text{Mixed}}(\theta)$ $=\underbrace{\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}}_{\text{on-policy}}\Bigg[\frac{1}{|\tau|}\sum_{t=1a}^{|\tau|}\operatorname{CLIP}\big(r_{t}(\theta),\,\hat{A}(\tau),\,\epsilon\big)\Bigg]$ (7)
$\quad+\underbrace{\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}^{\text{IEF}}}}_{\text{off-policy}}\Bigg[\frac{1}{|\tau|}\sum_{t=1}^{|\tau|}\operatorname{CLIP}\big(\hat{r}_{t}(\theta),\,\hat{A}(\tau),\,\epsilon\big)\Bigg],$
where $\hat{r}_{j,t}(\theta)=\frac{\pi_{\theta}(\tau_{j,t}\mid\tau_{j,<t})}{\pi^{\text{IEF}}_{\theta}(\tau_{j,t}\mid\tau_{j,<t})}$ is the expert-conditioned importance weight.
Unlike prior work Yan et al. (2025), which adopts a model-based off-policy scheme by relying on an additional advanced LRM $\pi_{\phi}$ to provide expert trajectories for the same training prompts, our mixed-policy GRPO with IEF operates as an input-conditioned off-policy method. Specifically, while all rollouts are sampled from the same policy $\pi_{\theta}$, in-context demonstrations alter the input conditioning and steer the policy away from its default output distribution. This conditioning discrepancy induces a behavior mismatch, under which expert-conditioned rollouts $\tau_{j}\sim\pi_{\theta}(x_{\text{exp}})$ are regarded as off-policy relative to standard on-policy samples $\tau_{i}\sim\pi_{\theta}(x)$.
4.2 Expert Region Reject Sampling
Building upon the expert-conditioned off-policy branch above, we further restrict updates to those trajectories that demonstrably improve model performance. We define an expert region as the subset of states where expert conditioning yields superior guidance, steering the policy beyond its native distribution. A rollout $\tau_{j}$ generated under expert conditioning is accepted into this region if its reward exceeds a predefined threshold $\delta$:
$\mathcal{E}_{\text{exp}}=\big\{(x_{\text{exp}},\tau_{j})\ \big|\ R(\tau_{j})\geq\delta\big\},$ (8)
where $\delta$ is set to $1.0$ by default.
To prevent low-quality expert-conditioned traces from biasing training, we define a reject sampling operator $\rho$ that selectively retains trajectories within the expert region. Formally, $\rho$ performs reject sampling by restricting the expectation to trajectories that fall within the expert region:
$\rho(\cdot)=\mathbb{E}_{\tau\sim\pi_{\theta}(\tau\mid\tau\in\mathcal{E}_{\text{exp}})}[g(\tau)],$ (9)
where $g(\tau)$ denotes the per-trajectory contribution to the objective. This filtering ensures that only high-reward expert-conditioned rollouts contribute to policy updates. The final objective of ICPO then becomes:
$\mathcal{J}_{\mathrm{ICPO}}(\theta)=\frac{1}{Z}\Big[\underbrace{\sum_{i=1}^{N_{\mathrm{on}}}\sum_{t=1}^{|\tau_{i}|}\mathrm{CLIP}(r_{i,t}(\theta),A_{i},\epsilon)}_{\text{on-policy objective}}$ (10)
$+\underbrace{\rho\left(\sum_{j=1}^{N_{\mathrm{off}}}\sum_{t=1}^{|\tau_{j}|}\mathrm{CLIP}(f(\hat{r}_{j,t}(\theta)),\hat{A}_{j},\epsilon)\right)}_{\text{off-policy objective}}\Big],$
where $Z$ normalizes over all valid tokens. The shaping function $f(\cdot)$ follows prior work and is defined as $f(x)=\frac{x}{x+\lambda}$, where $\lambda=0.01$ by default Yan et al. (2025). This shaping biases learning toward expert-induced improvements while encouraging exploration.
4.3 Reward Shaping with Annealed Expert Bonus
The verifiable reward function evaluates the model output by extracting the final answer and comparing it against the predefined ground-truth answer. It assigns a binary score based on whether the extracted answer matches the correct solution under a task-specific verifier. Formally,
$R(\tau)={\left\{\begin{array}[]{l l}{1}&{{\text{if}}\ \tau\ {\text{is correct}}}\\ {0}&{{\text{otherwise.}}}\end{array}\right.}$ (11)
This verifiable reward has been shown to reliably lead to successful scaling of RL training.
[…]
$R_{\text{shaped}}(\tau)=R(\tau)+\alpha\cdot\gamma(t),$ (12)
where $\gamma(t)=1-\frac{t}{T}$ denotes a linear decay scheduler over the training step $t$, and $\alpha$ denotes the bonus weight (set to $1.0$ in our experiments).
Algorithm 1 ICPO Training Procedure
0: Policy $\pi_{\theta}$, old policy $\pi_{\theta_{\text{old}}}$, expert data $\mathcal{D}$, batch size $B$, rollout size NN, few-shot count $k$, RS threshold $\delta$, step $t$, annealed bonus $\alpha\cdot\gamma(t)$
1:for each step do
2:  Sample prompts $\{x_{i}\}_{i=1}^{B}$
3:  for $i=1$ to $B$ do
4:   for $j=1$ to NN do
5:    $\tau_{i}^{j}\sim\pi_{\theta_{\text{old}}}(\cdot|x_{i})$
6:    Compute $R(\tau_{i}^{j})$
7:   end for
8:   Sample $k$ expert (q, a) pairs from $\mathcal{D}$ and form $x_{i}^{\text{exp}}$
9:   Generate $\tau_{i}^{\text{IEF}}\sim\pi_{\theta_{\text{old}}}^{\text{IEF}}(\cdot|x_{i}^{\text{exp}})$
10:   if $R(\tau_{i}^{\text{IEF}})\geq\delta$ and $\mathrm{correct}(\tau_{i}^{\text{IEF}})$ then
11:    Pick random $j$
12:    Replace $\tau_{i}^{j}\leftarrow\tau_{i}^{\text{IEF}}$
13:    if Enable reward shaping then
14:     $R(\tau_{i}^{j})\leftarrow R(\tau_{i}^{\text{IEF}})+\alpha\cdot\gamma(t)$
15:    end if
16:   end if
17:   Compute $\hat{A}_{i}$ using Eq. 6
18:  end for
19:  Compute mixed rollout loss $\mathcal{L}$ according to $\mathcal{J}_{\mathrm{ICPO}}(\theta)$
20:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
21:  $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
22:end for
5 Experimental Setup
Dataset.
We follow Yan et al. (2025) and adopt the OpenR1-Math-220k dataset as our main training corpus. Specifically, we use the filtered subset [^1] https://huggingface.co/datasets/Elliott/Openr1-Math-46k, which excludes generations exceeding 8192 tokens as well as those identified as incorrect by Math-Verify [^2] https://github.com/huggingface/Math-Verify. The resulting dataset contains approximately 45k verified reasoning prompts. For IEF, instead of using trajectories from advanced LRMs, we randomly sample demonstrations for each prompt from the MATH Hendrycks et al. (2021) training set, which contains 7.5k mathematical problems paired with high-quality solutions.
Implementation Details.
The complete set of hyperparameters used in our training is listed in Appendix B.1. For the main experiments, we use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base models and employ GRPO Shao et al. (2024) as our RL algorithm. We also include Qwen2.5-Math-7B to facilitate comparison with other related methods. We additionally evaluate ICPO on Qwen3-8B-Base and LLaMA-3.1-8B to examine its cross-model generalization. We generate a total of 8 rollout trajectories per prompt. For the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy rollout and 7 on-policy rollouts to ensure comparability. Evaluation details are provided in Appendix B.2.
Baseline Methods.
For direct comparison, we evaluate our ICPO against the vanilla GRPO baseline trained on the same subset of OpenR1-Math-220k. We further include $\text{GRPO}_{\text{ExtraRollouts}}$, which extends GRPO by using 16 on-policy rollouts to enhance exploration. To examine the effect of training data source, we also train a GRPO variant on expert-domain data, identical to our IEF demonstrations (MATH), denoted as $\text{GRPO}_{\text{ExpertDomain}}$. In addition, we compare with LUFFY Yan et al. (2025), which leverages trajectories generated by advanced LRMs as off-policy rollouts, along with other baselines detailed in Appendix B.3.
6 Experimental Results
6.1 Main Results
In-Distribution Benchmarks Out-of-Distribution Benchmarks
Method AIME 24/25 AMC23 MATH Minerva Olympiad Avg. (Impr.) ARC GPQA MMLU Avg. (Impr.)
Qwen3-1.7B 21.7 / 20.6 56.8 79.4 37.1 40.6 42.7 88.3 22.2 52.3 54.3
GRPO 28.4 / 22.5 66.7 83.6 40.8 48.2 48.4 88.3 34.3 54.4 59.0
$\text{GRPO}_{\text{ExtraRollouts}}$ 28.3 / 24.7 69.8 84.4 43.0 53.8 50.7 (+2.3) 88.9 30.8 55.0 58.2 (-0.8)
$\text{GRPO}_{\text{ExpertDomain}}$ 26.8 / 24.8 66.3 83.8 45.6 50.5 49.6 (+1.2) 88.8 32.8 55.5 59.0 (0.0)
ICPO (Ours) 31.3 / 26.3 70.4 86.8 44.1 56.4 52.5 (+4.1) 88.1 27.8 55.5 57.1 (-1.9)
[…]
Qwen3-8B 33.7 / 21.4 65.4 86.2 40.4 43.3 48.4 96.2 37.9 68.0 67.4
GRPO 54.8 / 38.5 83.8 91.0 50.7 62.4 63.5 95.8 51.0 72.0 72.9
$\text{GRPO}_{\text{ExtraRollouts}}$ 53.0 / 40.0 85.4 93.0 52.9 61.6 64.3 (+0.8) 95.7 53.0 72.9 73.9 (+1.0)
$\text{GRPO}_{\text{ExpertDomain}}$ 58.3 / 40.4 86.2 91.8 50.7 60.0 64.6 (+1.1) 92.2 51.0 70.2 71.1 (-1.8)
ICPO (Ours) 55.2 / 43.7 87.0 92.0 51.1 65.2 65.7 (+2.2) 95.5 55.1 72.3 74.3 (+1.4)
[…]
Method AIME24 AIME25 AMC23 MATH Minerva Olympiad Average
Qwen2.5-Math-7B Yang et al. (2025) 11.5 4.9 31.3 43.6 7.4 15.6 19.0
Previous RLVR Methods
SimpleRL-Zero* Zeng et al. (2025) 27.0 6.8 54.9 76.0 25.0 34.7 37.4
PRIME-Zero* Cui et al. (2025) 17.0 12.8 54.0 81.4 39.0 40.3 40.8
OpenReasoner-Zero* Hu et al. (2025b) 16.5 15.0 52.1 82.4 33.1 47.1 41.0
Oat-Zero* Liu et al. (2025b) 33.4 11.9 61.2 78.0 34.6 43.4 43.8
SFT and RL
SFT Yan et al. (2025) 23.8 22.9 61.8 82.2 37.9 42.1 45.1
[…]
Comparable Baseline Methods
ReLIFT* Ma et al. (2025) 28.2 22.9 64.8 85.0 37.1 54.9 48.8
LUFFY* Yan et al. (2025) 29.4 23.1 65.6 87.6 37.5 57.2 50.1
Prefix-RFT* Huang et al. (2025) 31.8 26.4 68.2 88.4 40.3 55.7 51.8
ICPO (Ours) 29.3 28.9 74.9 88.4 45.6 52.4 53.2
[…]
Table 2: Evaluation results for Qwen2.5-Math-7B. “*” indicate results reported from their original papers.
[…]
In-Distribution Evaluation.
[…]
These consistent gains across scales indicate that ICPO effectively steers the policy toward a more optimal output distribution. By leveraging both the challenging training set and high-quality expert demonstrations through IEF, ICPO forms a more informative training signal than either source alone—a synergistic effect that yields superior performance over both $\text{GRPO}_{\text{ExpertDomain}}$ and vanilla GRPO. Moreover, unlike ICPO, simply increasing the number of on-policy rollouts in $\text{GRPO}_{\text{ExtraRollouts}}$ offers only marginal improvement, revealing the inherent exploration limitations of vanilla GRPO and underscoring the importance of effective steering.
Out-of-Distribution Evaluation.
[…]
Comparison with Other Baselines.
[…]
(a) Expert Domain Reward
(b) In-Distribution Reward
(c) Training Reward
Figure 5: Reward curves of Qwen3-8B over training steps across test (Mean@2) and train sets.
6.2 Ablation Study
Effect of Each Component.
Variant MATH AIME24/25 AMC Mnrv. Avg.
Qwen3-1.7B
ICPO 86.8 31.3 / 26.3 70.4 44.1 51.8
- ERRS 85.6 32.2 / 25.9 66.8 42.3 50.6
- IEF (GRPO) 83.6 28.4 / 22.5 66.7 40.8 48.4
Qwen3-8B
ICPO 92.0 55.2 / 43.6 87.0 51.1 65.8
- ERRS 89.6 55.2 / 41.7 85.2 53.3 65.0
- IEF (GRPO) 91.0 54.8 / 38.5 83.8 50.7 63.8
RS = Reward Shaping, ERRS = Expert Region Reject Sampling, IEF = Implicit Expert Forcing.
Table 3: Ablation analysis by progressively removing components from ICPO.
Table 3 summarizes ablations by progressively removing components of ICPO. On Qwen3-8B, all components contribute positively: IEF yields the largest gain (+1.2) by injecting expert-conditioned guidance and enhancing exploration, while ERRS improves accuracy by filtering invalid expert-region trajectories (+0.8). Across both model sizes, removing any component consistently harms performance, demonstrating that all the modules are complementary and jointly essential for the effectiveness of ICPO.
Sources of Expert Guidance.
Variant A. E. R. MATH AIME24/25 AMC Mnrv. Avg.
Qwen3-1.7B
LUFFY 83.6 26.8 / 23.5 63.3 41.9 47.8
ICPO 85.6 32.2 / 25.9 66.8 42.3 50.6
ICPO 86.8 31.3 / 26.3 70.4 44.1 51.8
Qwen3-8B
LUFFY 91.0 53.1 / 37.0 85.3 52.2 63.7
ICPO 89.6 55.2 / 41.7 85.2 53.3 65.0
ICPO 92.0 55.2 / 43.6 87.0 51.1 65.8
A. = Advanced LRM Trajectory, E. = Existing Dataset, R. = Expert Region Reject Sampling (ERRS). ICPO = ICPO w/o ERRS.
Table 4: Comparison across expert guidance sources.
We compare ICPO with LUFFY Yan et al. (2025), which incorporates trajectories generated by advanced LRMs into off-policy GRPO, as shown in Table 4. ICPO, which removes ERRS and is thus directly comparable to LUFFY in its source of expert guidance, already surpasses LUFFY. This demonstrates that IEF can steer the model toward a better policy distribution by leveraging existing datasets as contextual guidance, eliminating the need for costly external LRM computation. Please refer to Appendix E for more ablation studies.
7 Analysis of Distribution Shift
To examine how ICPO alters the model’s predictive distribution, we conduct an in-depth analysis on the expert-domain dataset MATH-500 from both the instance- and token-level perspectives.
Instance-Level.
Figure 6 presents per-instance perplexity, where each point corresponds to an individual test case. Points lying below the diagonal $y=x$ indicate that the target method yields lower perplexity than GRPO on the same instance, suggesting a shift of the predictive distribution toward the expert domain. Simply increasing the number of on-policy rollouts yields no deviation from the GRPO baseline, indicating that rollout amplification alone does not alter the underlying policy distribution. In contrast, ICPO induces a clear shift toward expert-like reasoning patterns, and incorporating expert-bonus RS further strengthens it.
Token-Level.
[…]
Overall, the results demonstrate that ICPO effectively enables LRM to “think outside the policy.”
Figure 6: Perplexity as an instance-level measure of expert-domain distribution shift.
Figure 7: Token-rank shift as a token-level measure of expert-domain distribution shift.
8 Conclusion
We have presented ICPO, a unified RLVR framework that enhances reasoning without relying on external expert models. Leveraging the inherent ICL capability of LRMs, ICPO introduces mixed-policy GRPO with IEF, which constructs expert-conditioned rollouts from existing datasets, improving data utilization and expanding exploration beyond the current policy distribution. To ensure stable optimization, ICPO further integrates ERRS to eliminate noisy off-policy trajectories and adopts RS to facilitate a smooth transition from expert-guided imitation to autonomous optimization. Experiments show that ICPO consistently improves RL performance, highlighting its promise as a scalable and general post-training paradigm for LRMs.
Limitations
Despite the demonstrated improvements, our approach has several limitations that open avenues for future investigation.
Task Scope.
Our experiments are currently limited to mathematical reasoning tasks under the RLVR framework. Extending ICPO to other domains such as code generation would require additional task-specific adaptation.
Model Capability Assumption.
Our method assumes that the underlying model possesses basic ICL capability; without it, the implicit expert-conditioned rollouts may not yield meaningful supervision. Due to time constraints, we did not conduct experiments on models lacking ICL ability, such as training them to acquire this capability before applying GRPO.
Demonstration Sensitivity.
The model may be sensitive to the choice of ICL demonstrations. We experimented with various heuristic demonstration selection methods, as detailed in the Appendix C. The results show that randomly sampling demonstrations—both across batches and per prompt—yields the best performance. Consequently, we adopt random demonstration selection as the default strategy in our main experiments, without any deliberate curation. We hypothesize that carefully selecting demonstrations tailored to each problem could further improve performance, which we plan to explore in future work.
Expert Data Selection.
Our method eliminates the reliance on stronger external LRMs by leveraging the model’s own ICL capability to treat existing datasets as implicit experts, steering rollouts beyond the default GRPO distribution and enhancing exploration. Although ICPO improves the utility of available high-quality datasets through IEF, it still imposes several constraints on the expert data: (1) expert data must provide complete solutions; (2) the solutions must be of sufficiently high quality to offer effective guidance; and (3) using domain-specific expert data as demonstrations may bias the model toward that domain, potentially degrading its general-purpose capabilities, as detailed in Appendix D.3.
Despite these limitations, ICPO also opens promising directions. First, identifying high-quality, general-purpose expert datasets could further enhance model performance on both in-distribution and out-of-distribution tasks. Second, because IEF encourages the model to adopt the stylistic patterns of expert demonstrations, ICPO may naturally support controllable style transfer in reasoning tasks. We leave these directions to future work.
Ethics Statement
Use of AI Assistants.
We have employed ChatGPT as a writing assistant, primarily for polishing the text after the initial composition. We certify that any use of AI tools, including ChatGPT, was strictly limited to linguistic refinement such as improving grammar, clarity, and style. All substantive ideas, analyses, and arguments presented in this work originate from the authors or from properly cited prior research.
Computational Budget.
[…]
Reproducibility.
Our work is reproducible because we have provided our source code and implementation details.
Potential Risks.
To the best of our knowledge, there are no potential risks concerning our work.
Licenses.
The licenses of the scientific artifacts we use are shown in Table 5.
Category Artifact License
Model Qwen3 Models Apache-2.0
Qwen2.5 Models Apache-2.0
Framework LIMO MIT
VERL Apache-2.0
Math-Verify Apache-2.0
Dataset Skywork-OR1-RL-Data Apache-2.0
Openr1-Math-46k-8192 MIT
MathInstruct MIT
MATH MIT
MATH-500 MIT
Minerva MIT
OlympiadBench MIT
AMC23 N/A
AIME24 N/A
AIME25 N/A
Methods LUFFY N/A
Table 5: Licenses of scientific artifacts used in this work.
Appendix A Prompt Template
Here we provide the detailed prompt formats used for our experiments.
Prompt for Qwen3 # Prompt for RL and 0-shot Inference
<|im_start|>user
{QUESTION}
[…]
<|im_end|> <|im_start|>assistant
# Prompt for few-shot Inference
<|im_start|>user
[…]
{SYSTEM_PROMPT}
User:
Question: {QUESTION}
Answer:
Assistant: <think>
# Prompt for few-shot Inference
{SYSTEM_PROMPT}
User:
Question: {QUESTION}
Answer: {ANSWER} $n \n$
Question: {QUESTION}
Answer:
Assistant: <think>
Figure 8: Prompt formats used for RL training, zero-shot inference, and few-shot inference.
Appendix B Experimental Details
B.1 Training Settings
For RL finetuning, we use the widely adopted GRPO algorithm built on VERL Sheng et al. (2025) framework. The full hyperparameters used in our training are listed in Table 6. Specifically, for rollout generation, we use a temperature of 1.0, and rewards are computed using Math-Verify. All models are trained for $T=400$ optimization steps, and we report results using the final checkpoint.
Hyper-parameter Value
Learning Rate 1e-6
Totel Steps 400
Batch Size 128
Mini Batch Size 64
KL Loss Coefficient 0.0
Clip Ratio 0.2
Temperature 1.0
Total Number of Rollouts 8
Maximum Prompt Length 4096
Maximum Response Length 8192
Table 6: Full hyper-parameters for training.
B.2 Evaluation Settings
We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al. (2024), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), and MATH-500 Hendrycks et al. (2021). For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1. As a complementary analysis, we also report Pass@32 results in Appendix D.2.
To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al. (2018), GPQA-Diamond Rein et al. (2023), and MMLU-Pro Wang et al. (2024), with multiple-choice options shuffled to prevent contamination. All evaluations are conducted using the LIMO framework Ye et al. (2025). During inference, we follow Yan et al. (2025) and set the generation temperature to 0.6. For ICL evaluation, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set using 5 different random seeds, and report the average performance across them.
B.3 Baseline Methods
We benchmark ICPO against the following baselines using Qwen2.5-Math-7B Yang et al. (2024) as the base model. We compare our method with three lines of work, which is previous RLVR methods, SFT and RL baseline, and methods combining SFT and RL that is comparable with our ICPO.
Previous RLVR Methods.
(1) SimpleRL-Zero Zeng et al. (2025), which applies GRPO to approximately 24k mathematical samples from GSM8K Cobbe et al. (2021) and MATH dataset Hendrycks et al. (2021); (2) PRIME-Zero Cui et al. (2025), which conducts policy rollouts on 150k NuminaMath queries with implicit process rewards; (3) OpenReasoner-Zero Hu et al. (2025b), a PPO-based approach trained on 129k multi-source samples; and (4) Oat-Zero Liu et al. (2025b), which removes the standard deviation in GRPO advantage calculation, and is trained on the MATH dataset.
SFT and RL.
[…]
Comparable Baseline Methods.
For methods that integrate supervised fine-tuning (SFT) with reinforcement learning (RL), we evaluate three representative approaches: (1) LUFFY Yan et al. (2025), a mixed-policy GRPO method that incorporates advanced LRM reasoning trajectories as off-policy rollouts; (2) ReLIFT Ma et al. (2025), which alternates standard RLVR updates with targeted supervised fine-tuning on the hardest online-collected questions, enabling the model to acquire new reasoning skills that pure RL cannot provide; (3) Prefix-RFT Huang et al. (2025), which samples a prefix from advanced LRM reasoning trajectories and reinforces the model’s continuation, mixing these hybrid sequences with online rollouts under the RFT objective to unify SFT-style imitation and RFT-style exploration.
All these methods rely on advanced LRM-generated reasoning trajectories from the training set as demonstrations, whereas ICPO leverages few-shot demonstrations retrieved from an existing external dataset—independent of the training set—and uses them purely as in-context prompts to steer the rollout distribution on-policy.
B.4 Data Statistics
Figure 9 presents the length distribution of the demonstrations (question + answer) and the training prompts. We observe that the demonstrations are relatively short, ensuring that they do not cause input truncation or the loss of important information. Compared to approaches that append the model’s generated reasoning traces as additional context, our overall input length remains much shorter and thus more efficient.
Figure 9: Length distribution of expert demonstrations (left) and training prompts (right).
Appendix C Demonstration Selection Strategies
Figure 10: Rewards on the MATH-500 dataset for different heuristic demonstration-selection strategies using Qwen3-1.7B. The Hard, Short, and Random32 demonstration strategies achieve comparable performance and all outperform the GRPO baseline.
Figure 11: Comparison of ICPO training using different demonstration pool sizes on Qwen3-8B. The left figure reports the reward and training dynamics on the MATH-500 benchmark, while the right figure presents the number of non-pass cases on the training set.
Effects of Heuristic Demonstration Selection Strategies.
We investigate several heuristic strategies for selecting demonstrations within ICPO. Experiments are conducted on the Skywork-OR1-RL-Data He et al. (2025) dataset using Qwen3-1.7B, with all demonstrations sourced from the training split of MATH Hendrycks et al. (2021) as our main setup. We evaluate the following heuristic selection methods:
  • Difficulty-based selection. We sample each MATH training instance three times: examples solved correctly in all three attempts are labeled Easy, those answered incorrectly three times are labeled Hard, and the remaining examples are categorized as Medium. During training, each prompt is matched with a randomly selected demonstration from its corresponding difficulty group.
  • Length-based selection. We partition both the demonstration pool and the training set into two equal-sized buckets based on response length. For each training sample, a demonstration is randomly selected from the bucket with the same length range.
  • Subject-based selection. We group both the demonstration data and the training data by their problem category (subject). Each training instance is paired with a randomly selected example from the same subject as its 1-shot demonstration.
  • Random. We randomly sample 32 examples from the entire demonstration pool and, for each training instance at every step, randomly select one of these 32 as the demonstration.
The results are shown in Figure 10. Among all heuristic strategies, the Hard, Short, and Random32 variants achieve comparable performance, and all surpass the GRPO baseline. Given that difficulty-based and length-based methods require additional sampling or preprocessing cost, we adopt the simplest and most cost-efficient option, namely Random demonstration selection, as our default strategy.
Effects of Demonstration Pool Size.
We further investigate how the size of the demonstration pool available to each batch affects ICPO training using Qwen3-8B, with the results shown in Figure 11. We compare two configurations. The first sets the demo pool size to 10 (denoted as ICPO-10/7500): for each batch, we randomly sample 10 demonstrations, and each instance within the batch randomly selects one of these as its 1-shot demonstration. This design aims to prevent excessive steering directions within a batch, which may destabilize training. The second configuration sets the demonstration pool size equal to the full demonstration set (denoted as ICPO-7500/7500), where each training instance samples its 1-shot demonstration from the entire pool of 7,500 examples, thereby increasing exploration during training.
As shown in the results, ICPO-10/7500 yields more stable and stronger early-stage performance but restricts exploration in later stages. In contrast, ICPO-7500/7500 demonstrates greater potential in the later phase, achieving higher rewards on the expert domain (MATH-500) and resolving more non-pass cases. Based on these observations, we adopt random selection with a full demonstration pool size of 7,500 as the default configuration for our main experiments.
Appendix D Results and Analysis
D.1 Training Dynamics
Figure 14 illustrates the training dynamics. For smaller models, ICPO maintains a higher policy entropy than GRPO, reflecting a broader policy support and increased exploration during training.
For larger models, their stronger ICL capability makes IEF more effective. This leads to higher policy entropy and KL divergence, while we retain the KL loss to provide a valid constraint on policy updates. As the stronger exploration, correspondingly, the generated responses become longer.
Overall, these observations indicate that ICPO facilitates stronger exploration compared to GRPO.
(a) Expert Domain Reward
(b) In-Distribution Reward
(c) Training Reward
Figure 12: Reward curves of Qwen3-1.7B over training steps across test and train sets.
(a) Training dynamics of Qwen3-1.7B.
(b) Training dynamics of Qwen3-8B.
Figure 14: Training dynamics of ICPO.
D.2 Pass@32 Performance
We report Pass@32 results to further analyze the exploration behavior of our ICPO. As shown in Table 7, simply increasing the number of rollouts ($\text{GRPO}_{\text{ExtraRollouts}}$) or directly training on expert-domain data ($\text{GRPO}_{\text{ExpertDomain}}$) yields only limited or inconsistent improvements in Pass@32 across benchmarks. In contrast, ICPO achieves higher scores across both Qwen3-8B-Base and Qwen3-8B, indicating that IEF and RS during RL training effectively guides exploration toward expert-aligned regions of the solution space, enlarging the set of reachable correct solutions rather than merely sharpening the policy around a single mode. This behavior aligns with our motivation of leveraging LRM’s inherent ICL ability to enhance exploration without relying on advanced LRMs.
Method AIME24 AIME25 AMC23 MATH-500 Minerva Olympiad Average
Qwen3-8B-Base 46.7 40.0 85.0 94.2 65.4 68.3 66.6
GRPO 53.3 43.3 90.0 96.2 58.8 72.7 69.1
$\text{GRPO}_{\text{ExtraRollouts}}$ 50.0 46.7 92.5 95.4 64.3 73.8 70.4
$\text{GRPO}_{\text{ExpertDomain}}$ 46.7 26.7 90.0 92.0 61.0 64.2 63.4
ICPO (Ours) 60.0 60.0 95.0 95.8 63.6 73.0 74.6
[…]
Qwen3-8B 56.7 36.7 85.0 93.2 57.7 57.5 64.5
GRPO 80.0 60.0 95.0 97.0 65.1 79.0 79.3
$\text{GRPO}_{\text{ExtraRollouts}}$ 80.0 60.0 95.0 97.0 62.9 79.4 79.0
$\text{GRPO}_{\text{ExpertDomain}}$ 83.3 66.7 92.5 97.8 64.7 77.6 80.4
ICPO (Ours) 86.7 66.7 95.0 96.8 65.4 75.0 80.9
[…]
Table 7: Pass@32 results for Qwen3-8B-Base and Qwen3-8B.
D.3 OOD Performance
We evaluate ICPO and the corresponding baselines on out-of-distribution (OOD) benchmarks using Qwen2.5 models, as shown in Table 2 and Table 8. Although Qwen2.5-Math-7B performs reasonably well on in-distribution datasets, it fails to outperform other baselines on OOD evaluations. In contrast, stronger base models such as Qwen3 achieve consistently superior performance on both in-distribution and OOD settings (see Table 1).
We hypothesize that this discrepancy stems primarily from differences in the base models’ inherent ICL capabilities. Models with stronger in-context learning ability exhibit more robust generalization, and consequently achieve better overall performance across both in-distribution and OOD tasks.
Method ARC-C GPQA-D MMLU-Pro Average
Qwen2.5-Math-7B Yang et al. (2025) 18.2 11.1 16.9 15.4
Previous RLVR Methods
SimpleRL-Zero* Zeng et al. (2025) 30.2 23.2 34.5 29.3
PRIME-Zero* Cui et al. (2025) 73.3 18.2 32.7 41.4
OpenReasoner-Zero* Hu et al. (2025b) 66.2 29.8 58.7 51.6
Oat-Zero* Liu et al. (2025b) 70.1 23.7 41.7 45.2
SFT and RL
SFT Yan et al. (2025) 74.7 28.3 44.4 49.1
RL Shao et al. (2024) 75.9 32.8 42.6 50.4
SFT+RL Shao et al. (2024) 79.2 37.9 49.6 55.5
Comparable Baseline Methods
ReLIFT Ma et al. (2025) 74.9 40.9 51.9 55.9
LUFFY Yan et al. (2025) 80.5 39.9 53.0 57.8
Prefix-RFT Huang et al. (2025) 84.0 39.1 52.1 58.4
ICPO (Ours) 74.0 34.3 46.5 51.6
[…]
Table 8: Evaluation results for Qwen2.5-Math-7B. Methods marked with “*” indicate results reported from their original papers.
D.4 Generalizability Across Models
Method AIME24 AIME25 AMC23 MATH-500 Minerva Olympiad Average
Qwen3-8B-Base 10.8 10.4 47.6 67.2 32.0 34.4 33.7
GRPO 22.5 21.4 71.8 84.6 43.8 50.1 49.0
$\text{GRPO}_{\text{ExtraRollouts}}$ 26.9 20.2 67.4 85.8 46.7 51.6 49.8
$\text{GRPO}_{\text{ExpertDomain}}$ 22.9 12.8 60.9 80.0 47.8 42.8 44.5
LUFFY 24.6 22.1 68.2 87.2 47.4 52.3 50.3
ICPO 26.8 21.6 69.2 88.6 46.3 53.8 51.0
[…]
Table 9: Results on mathematical reasoning benchmarks for Qwen3-8B-Base.
Method AIME24 AIME25 AMC23 MATH-500 Minerva Olympiad Average
LLaMA-3.1-8B-Instruct 4.6 0.2 20.9 45.4 22.8 15.7 18.3
SFT 0.5 0.1 5.4 20.2 4.0 5.3 5.9
GRPO 3.1 1.0 10.8 28.2 17.6 16.2 12.8
LUFFY 1.9 0.1 13.5 39.0 15.1 9.6 13.2
ICPO 3.1 0.2 16.6 33.4 21.0 9.0 13.9
[…]
Table 10: Results on mathematical reasoning benchmarks for LLaMA-3.1-8B Grattafiori et al. (2024). Methods marked with “*” indicate results reported from Yan et al. (2025).
To evaluate the generalizability of our method across various model architectures, we extend ICPO to the base model Qwen3-8B-Base and LLaMA-3.1-8B. As shown in Table 9 and Table 10, ICPO variants surpass SFT, RL baseline and LUFFY Yan et al. (2025) on six mathematical benchmarks, underscoring its reasoning performance.
Appendix E Ablation Study
E.1 Ablation on Expert Guidance
Expert-Domain Data.
Variant MATH AIME24/25 AMC Mnrv. Avg.
GRPO 91.0 54.8 / 38.5 83.8 50.7 63.8
MATH Hendrycks et al. (2021)
ICPO 92.0 55.2 / 43.6 87.0 51.1 65.8
[…]
MathInstruct Yue et al. (2023)
ICPO 90.8 56.7 / 40.7 83.8 53.7 65.1
[…]
Table 11: Ablation analysis across different expert-domain data on Qwen3-8B.
We study the generalizability of ICPO under different sources of expert-domain data. In addition to our default setting, we adopt MathInstruct [^3] https://huggingface.co/datasets/TIGER-Lab/MathInstruct Yue et al. (2023) as an alternative expert demonstration corpus for ICPO. MathInstruct contains approximately 262k instruction–solution pairs. To enable reliable extraction of ground-truth answers for RLVR, we select a subset of 118k examples whose solutions explicitly include the phrase “The answer is”. As shown in Table 11, ICPO and its variants guided by different expert-domain data sources consistently outperform the GRPO baseline, demonstrating the robustness of our approach to the choice of expert data. Notably, although all evaluations are conducted on in-distribution benchmarks, expert data drawn from different domains exhibits varying strengths across benchmarks. This observation suggests that more carefully curated or benchmark-aligned demonstrations may further improve generalization performance across diverse tasks.
E.2 Ablation on Difficulty Levels
Train Data.
Method MATH AIME24/25 AMC Mnrv. Avg.
OpenR1-Math-220k
GRPO 83.6 28.4 / 22.5 66.7 40.8 48.4
ICPO 86.8 31.3 / 26.3 70.4 44.1 51.8
Skywork-OR1-RL-Data
GRPO 83.0 25.1 / 22.2 66.9 42.3 47.9
ICPO 86.0 26.8 / 24.1 69.6 42.6 49.8
Table 12: Comparison of ICPO performance on Qwen3-1.7B under two training regimes.
We additionally train on a simpler dataset, Skywork-OR1-RL-Data [^4] https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al. (2025), to verify the generalization ability of our IEF under different reasoning conditions. This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison. As shown in Table 12, results verify that ICPO is consistently beneficial across training datasets of different difficulty levels. On the simpler Skywork dataset, ICPO improves GRPO by +1.9 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.4 average points. This demonstrates that ICPO not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
Prompt Groups.
Variant MATH AIME24/25 AMC Mnrv. Avg.
GRPO 83.0 25.1 / 22.2 66.9 42.3 47.9
+IEF on Some-Pass 82.0 26.9 / 26.0 67.8 43.4 49.2
+IEF on Non-Pass 82.4 25.42 / 23.33 67.19 42.28 48.12
ICPO 86.0 26.8 / 24.1 69.6 42.7 49.8
Table 13: Ablation study of IEF on Qwen3-1.7B using the Skywork dataset, grouped by rollout success rates.
Table 13 further analyzes the effectiveness of ICPO by grouping prompts according to their initial rollout success rates. We denote prompts with a success rate of 0.0 as Non-Pass, and those with success rates between 0.0 and 1.0 as Some-Pass. ICPO applies implicit expert forcing (IEF) across all prompt groups.
We observe that expert guidance consistently improves performance for both Some-Pass and Non-Pass prompts. Moreover, the impact of IEF differs across prompt types and exhibits complementary benefits. For Some-Pass prompts, IEF enriches solution diversity while remaining within expert-aligned regions, thereby mitigating premature policy specialization. For Non-Pass prompts, IEF provides reliable expert guidance that enables the policy to solve prompts that cannot be handled by vanilla GRPO.
Taken together, these results demonstrate that ICPO enhances exploration for Some-Pass cases while enabling effective learning on challenging Non-Pass prompts.
Appendix F Token Rank Shift Analysis.
To quantify how different training methods alter model behavior at a fine-grained level, we follow Lin et al. (2023) and adopt token-rank shift as a token-level diagnostic metric. The core idea is to measure how unlikely the generated tokens of an aligned model are under the base model distribution.
Given a generated trajectory produced by an aligned policy, we replay the same trajectory under the base model and examine, at each generation step, how the aligned token is ranked according to the base model’s conditional distribution. Let $x_{<t}$ denote the prefix up to position t1t-1, and let $y_{t}$ be the token generated by the aligned model at step $t$. We compute the base model logits $\ell^{\text{base}}(\cdot\mid x_{<t})$ and define the base-rank of $y_{t}$ as:
$\mathrm{rank}(y_{t})\;=\;1+\left|\left\{v\;\middle|\;\ell^{\text{base}}(v\mid x_{<t})>\ell^{\text{base}}(y_{t}\mid x_{<t})\right\}\right|.$ (13)
A rank of $1$ indicates that the aligned token coincides with the base model’s top-1 prediction, while larger ranks indicate increasing deviation from the base policy’s high-probability region.
For interpretability, we further categorize token ranks into three regimes: (1) Unshifted: $\mathrm{rank}(y_{t})=1$; (2) Marginally shifted: $1<\mathrm{rank}(y_{t})\leq k$; (3) Shifted: $\mathrm{rank}(y_{t})>k$, where $k$ is a small threshold (e.g., $k=3$).
We define the shifted ratio as the fraction of tokens whose base-rank exceeds $k=3$, which serves as a proxy for how frequently the aligned policy samples tokens that would be unlikely under the base model distribution.
Due to the varying lengths of reasoning trajectories, analyzing token rank shift using absolute token positions can be misleading. Instead, we normalize token positions by trajectory length and compute statistics over length percentiles. In Figure 7, we discretize the normalized positions into 100 bins and aggregate token shift metrics within each bin. By aggregating these statistics across token positions, token rank shift enables a fine-grained analysis of when and to what extent the aligned policy departs from the base policy during generation.

### v3

@@
 Results demonstrate that ICPO consistently enhances RLVR performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs.
-Our code is available at https://anonymous.4open.science/r/ICPO.
+Our code is available at https://github.com/Celine-hxy/ICPO.
 Think Outside the Policy: In-Context Steered Policy Optimization
-Hsiu-Yuan Huang1,2thanks:   Equal contribution.
-Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3thanks:   Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1, [^2] footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉  huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
+Hsiu-Yuan Huang1,2,3thanks:   Equal contribution.
+Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3thanks:   Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1, [^2] footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University  3LLM Department, Tencent Correspondence: huang.hsiuyuan@stu.pku.edu.cn wuyf@pku.edu.cn
 1 Introduction
@@
 (2024); Hojel et al.
+(2024); Cai et al.
 (2025), and $\mathcal{F}(\cdot\,;\vartheta)$ represents the task-specific reasoning function that applies the $\vartheta$ to generate the prediction for query $q$.
@@
 (2025) and use 1 off-policy rollout and 7 on-policy rollouts to ensure comparability.
-Evaluation details are provided in Appendix B.2.
+Evaluation Settings.
+We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al.
+(2024), Minerva Lewkowycz et al.
+(2022), Olympiad He et al.
+(2024), and MATH-500 Hendrycks et al.
+(2021).
+To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al.
+(2018), GPQA-Diamond Rein et al.
+(2023), and MMLU-Pro Wang et al.
+(2024), with multiple-choice options shuffled to prevent contamination.
+For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1.
+More evaluation details are provided in Appendix B.2.
 Baseline Methods.
@@
 (c) Training Reward
-Figure 5: Reward curves of Qwen3-8B over training steps across test (Mean@2) and train sets.
+Figure 5: Reward curves over training steps across test (Mean@2) and train sets.
+Figure 6: Training dynamics of Qwen3-1.7B.
 6.2 Ablation Study
@@
 Across both model sizes, removing any component consistently harms performance, demonstrating that all the modules are complementary and jointly essential for the effectiveness of ICPO.
+Selection of Expert Data.
+Method MATH AIME24/25 AMC Mnrv. Avg.
+Qwen3-1.7B
+GRPO 83.6 28.4 / 22.5 66.7 40.8 48.4
+ICPO (PoT) 87.0 31.5 / 27.0 69.6 42.3 51.5
+ICPO (CoT) 86.8 31.3 / 26.3 70.4 44.1 51.8
+Qwen3-8B
+GRPO 91.0 54.8 / 38.5 83.8 50.7 63.8
+ICPO (PoT) 90.8 56.7 / 40.7 83.8 53.7 65.1
+ICPO (CoT) 92.0 55.2 / 43.6 87.0 51.1 65.8
+Table 4: Ablation on expert data selection, where CoT refers to mathematical data (MATH) and PoT refers to code-oriented OOD reasoning data.
+To demonstrate the robustness of ICPO to the choice of expert data, we replace the in-domain expert data with cross-domain Program-of-Thought (PoT) data Yue et al.
+(2023) and conduct ICPO training under this setting.
+Details of data processing are provided in Appendix E.1.
+As shown in Table 4, ICPO consistently outperforms GRPO when using either CoT or PoT data as demonstrations for implicit expert guidance.
+These results suggest that ICPO is robust to the choice of expert data and highlight its potential to generalize beyond mathematical reasoning to cross-domain settings.
+More importantly, by changing the expert data, ICPO functions as a plug-and-play framework that flexibly reshapes the target policy distribution, enabling controllable steering of model behavior.
 Sources of Expert Guidance.
@@
 Please refer to Appendix E for more ablation studies.
+Role of Base Model ICL Capability.
+Method MATH AIME24/25 AMC Mnrv. Avg.
+Pass@1
+GRPO 84.6 22.5 / 21.4 71.8 43.8 48.8
+$\text{GRPO}_{\text{ExtraRollouts}}$ 85.8 26.9 / 20.2 67.4 46.7 49.4
+$\text{GRPO}_{\text{ExpertDomain}}$ 80.0 22.9 / 12.8 60.9 47.8 44.9
+ICPO (Ours) 88.6 26.8 / 21.6 69.2 46.3 50.5
+Pass@32
+GRPO 96.2 53.3 / 43.3 90.0 58.8 68.3
+$\text{GRPO}_{\text{ExtraRollouts}}$ 95.4 50.0 / 46.7 92.5 64.3 69.8
+$\text{GRPO}_{\text{ExpertDomain}}$ 92.0 46.7 / 26.7 90.0 61.0 63.3
+ICPO (Ours) 95.8 60.0 / 60.0 95.0 63.6 74.9
+Table 6: Pass@1 and Pass@32 results for Qwen3-8B-Base, a model with no ICL capability.
+To examine whether ICPO relies on the assumption that the base model possesses basic ICL capability, we conduct additional experiments on Qwen3-8B-Base, a model with no ICL capability.
+The results are summarized in Table 6.
+The observed Pass@1 improvement demonstrates that ICPO remains effective even when the base model lacks ICL ability.
+Meanwhile, the substantial Pass@32 improvement indicates that ICPO expands the exploration space rather than merely sharpening existing modes, encouraging the model to “think outside the policy.” These results show that ICPO is not predicated on the ICL capabilities of the base model Instead, ICL acts as a scaling amplifier: stronger ICL ability leads to larger improvements, yet weakly aligned models can also benefit from implicit expert forcing during RL optimization.
 7 Analysis of Distribution Shift
@@
 Limitations
-Despite the demonstrated improvements, our approach has several limitations that open avenues for future investigation.
-Task Scope.
-Our experiments are currently limited to mathematical reasoning tasks under the RLVR framework.
-Extending ICPO to other domains such as code generation would require additional task-specific adaptation.
-Model Capability Assumption.
-Our method assumes that the underlying model possesses basic ICL capability; without it, the implicit expert-conditioned rollouts may not yield meaningful supervision.
-Due to time constraints, we did not conduct experiments on models lacking ICL ability, such as training them to acquire this capability before applying GRPO.
-Demonstration Sensitivity.
-The model may be sensitive to the choice of ICL demonstrations.
-We experimented with various heuristic demonstration selection methods, as detailed in the Appendix C.
-The results show that randomly sampling demonstrations—both across batches and per prompt—yields the best performance.
-Consequently, we adopt random demonstration selection as the default strategy in our main experiments, without any deliberate curation.
-We hypothesize that carefully selecting demonstrations tailored to each problem could further improve performance, which we plan to explore in future work.
-Expert Data Selection.
-Our method eliminates the reliance on stronger external LRMs by leveraging the model’s own ICL capability to treat existing datasets as implicit experts, steering rollouts beyond the default GRPO distribution and enhancing exploration.
-Although ICPO improves the utility of available high-quality datasets through IEF, it still imposes several constraints on the expert data: (1) expert data must provide complete solutions; (2) the solutions must be of sufficiently high quality to offer effective guidance; and (3) using domain-specific expert data as demonstrations may bias the model toward that domain, potentially degrading its general-purpose capabilities, as detailed in Appendix D.3.
-Despite these limitations, ICPO also opens promising directions.
-First, identifying high-quality, general-purpose expert datasets could further enhance model performance on both in-distribution and out-of-distribution tasks.
-Second, because IEF encourages the model to adopt the stylistic patterns of expert demonstrations, ICPO may naturally support controllable style transfer in reasoning tasks.
-We leave these directions to future work.
+Our experiments focus primarily on mathematical reasoning and OOD reasoning benchmarks under the RLVR framework, including multi-domain knowledge reasoning (MMLU), open-domain QA (GPQA), and scientific reasoning tasks (ARC).
+Extending ICPO to other domains, such as code generation, may require task-specific adaptations, especially in reward design and evaluation.
+We leave a systematic investigation of these broader applications to future work.
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China (62076008).
 Ethics Statement
@@
 B.2 Evaluation Settings
-We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al.
-(2024), Minerva Lewkowycz et al.
-(2022), Olympiad He et al.
-(2024), and MATH-500 Hendrycks et al.
-(2021).
 For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1.
 As a complementary analysis, we also report Pass@32 results in Appendix D.1.
-To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al.
-(2018), GPQA-Diamond Rein et al.
-(2023), and MMLU-Pro Wang et al.
-(2024), with multiple-choice options shuffled to prevent contamination.
 All evaluations are conducted using the LIMO framework Ye et al.
@@
 Appendix D Results and Analysis
-D.1 Training Dynamics
-Figure 14 illustrates the training dynamics. For smaller models, ICPO maintains a higher policy entropy than GRPO, reflecting a broader policy support and increased exploration during training.
-For larger models, their stronger ICL capability makes IEF more effective.
-This leads to higher policy entropy and KL divergence, while we retain the KL loss to provide a valid constraint on policy updates.
-As the stronger exploration, correspondingly, the generated responses become longer.
-Overall, these observations indicate that ICPO facilitates stronger exploration compared to GRPO.
 (a) Expert Domain Reward
@@
 Figure 13: Reward curves of Qwen3-1.7B over training steps across test and train sets.
-(a) Training dynamics of Qwen3-1.7B.
-(b) Training dynamics of Qwen3-8B.
-Figure 14: Training dynamics of ICPO.
-D.2 Pass@32 Performance
+D.1 Pass@32 Performance
 We report Pass@32 results to further analyze the exploration behavior of our ICPO.
@@
 Table 9: Pass@32 results for Qwen3-8B-Base and Qwen3-8B.
-D.3 OOD Performance
+D.2 OOD Performance
 We evaluate ICPO and the corresponding baselines on out-of-distribution (OOD) benchmarks using Qwen2.5 models, as shown in Table 2 and Table 10.
@@
 Table 10: Evaluation results for Qwen2.5-Math-7B. Methods marked with “*” indicate results reported from their original papers.
-D.4 Generalizability Across Models
+D.3 Generalizability Across Models
 Method AIME24 AIME25 AMC23 MATH-500 Minerva Olympiad Average
@@
 Expert-Domain Data.
-Variant MATH AIME24/25 AMC Mnrv. Avg.
-GRPO 91.0 54.8 / 38.5 83.8 50.7 63.8
-MATH Hendrycks et al. (2021)
-ICPO 92.0 55.2 / 43.6 87.0 51.1 65.8
-[…]
-MathInstruct Yue et al. (2023)
-ICPO 90.8 56.7 / 40.7 83.8 53.7 65.1
-[…]
-Table 11: Ablation analysis across different expert-domain data on Qwen3-8B.
 We study the generalizability of ICPO under different sources of expert-domain data.
@@
 To enable reliable extraction of ground-truth answers for RLVR, we select a subset of 118k examples whose solutions explicitly include the phrase “The answer is”.
-As shown in Table 11, ICPO and its variants guided by different expert-domain data sources consistently outperform the GRPO baseline, demonstrating the robustness of our approach to the choice of expert data.
-Notably, although all evaluations are conducted on in-distribution benchmarks, expert data drawn from different domains exhibits varying strengths across benchmarks.
-This observation suggests that more carefully curated or benchmark-aligned demonstrations may further improve generalization performance across diverse tasks.
+As shown in Table 4, ICPO consistently outperforms GRPO when using either CoT or PoT data as demonstrations for implicit expert guidance.
+These results suggest that ICPO is robust to the choice of expert data, and highlight its potential to generalize beyond mathematical reasoning to cross-domain settings, such as code-oriented tasks.
 E.2 Ablation on Difficulty Levels
 Train Data.
+We additionally train on a simpler dataset, Skywork-OR1-RL-Data [^4] https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al.
+(2025), to verify the generalization ability of our IEF under different reasoning conditions.
+This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison.
+As shown in Table 13, results verify that ICPO is consistently beneficial across training datasets of different difficulty levels.
+On the simpler Skywork dataset, ICPO improves GRPO by +1.9 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.4 average points.
+This demonstrates that ICPO not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
 Method MATH AIME24/25 AMC Mnrv. Avg.
@@
 Table 13: Comparison of ICPO performance on Qwen3-1.7B under two training regimes.
-We additionally train on a simpler dataset, Skywork-OR1-RL-Data [^4] https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al.
-(2025), to verify the generalization ability of our IEF under different reasoning conditions.
-This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison.
-As shown in Table 12, results verify that ICPO is consistently beneficial across training datasets of different difficulty levels.
-On the simpler Skywork dataset, ICPO improves GRPO by +1.9 average points, while on the more challenging OpenR1-Math corpus, the gain further increases to +3.4 average points.
-This demonstrates that ICPO not only enhances learning on complex reasoning traces but also generalizes effectively to settings with weaker supervision.
 Prompt Groups.
-Variant MATH AIME24/25 AMC Mnrv. Avg.
-GRPO 83.0 25.1 / 22.2 66.9 42.3 47.9
-+IEF on Some-Pass 82.0 26.9 / 26.0 67.8 43.4 49.2
-+IEF on Non-Pass 82.4 25.42 / 23.33 67.19 42.28 48.12
-ICPO 86.0 26.8 / 24.1 69.6 42.7 49.8
-Table 13: Ablation study of IEF on Qwen3-1.7B using the Skywork dataset, grouped by rollout success rates.
 Table 14 further analyzes the effectiveness of ICPO by grouping prompts according to their initial rollout success rates.
@@
 Taken together, these results demonstrate that ICPO enhances exploration for Some-Pass cases while enabling effective learning on challenging Non-Pass prompts.
+Variant MATH AIME24/25 AMC Mnrv. Avg.
+GRPO 83.0 25.1 / 22.2 66.9 42.3 47.9
++IEF on Some-Pass 82.0 26.9 / 26.0 67.8 43.4 49.2
++IEF on Non-Pass 82.4 25.4 / 23.3 67.2 42.3 48.1
+ICPO 86.0 26.8 / 24.1 69.6 42.7 49.8
+Table 14: Ablation study of IEF on Qwen3-1.7B using the Skywork dataset, grouped by rollout success rates.
 Appendix F Token Rank Shift Analysis.